OpenStudy (anonymous):

8 3/4 divided by 2 1/6=?

OpenStudy (anonymous):

it needs to be a fraction in its simplest form

OpenStudy (anonymous):

I always hated fractions..lol ok..here we go....turn your mixed numbers to improper fractions 8 3/4 = (8 * 4 + 3)/4 = 35/4 2 1/6 = (2 * 6 + 1) / 6 = 13/6 so now your problem is : 35/4 divided by 13/6 when dividing fractions, you " flip" what you are dividing by, then multiply. 35/4 * 6/13 -- notice I flipped what I was dividing by and not the other number. Now we multiply 210/52 and if you need it back in mixed number, it is 4 1/26

OpenStudy (anonymous):

or this might be what they want.. 210/52 = 105/26 I do not know if you need the answer as an improper fraction 105/26 or as a mixed number 4 1 2/6

OpenStudy (anonymous):

oops..my bad...mixed number is 4 1/26

OpenStudy (anonymous):

as a mixed number

OpenStudy (anonymous):

4 1/26 ( 4 and 1/26) is your mixed number

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

YW :)

OpenStudy (anonymous):

:)